OpenStudy (anonymous):

A metalworker has a metal alloy that is 20% copper and another alloy that is 55% copper. How many kilograms of each alloy should the metalworker combine to create 100kg of a 48% copper alloy? The metal worker should use how many kilograms of the metal alloy that is 20% copper and how many kilograms of the metal alloy that is 55% copper?

OpenStudy (anonymous):

1 kg of alloy(1) 20% coppre 1 kg of alloy (2) 55% copper copper % of these 2 kg = (20+55)/2=37.5%

OpenStudy (anonymous):

this is a hint... (alloy (1) mass * alloy(1) copper%)+(alloy(2)mass*(alloy(2) copper%)/total mass=48 copper %

OpenStudy (anonymous):

mass (1)+mass(2)=100

OpenStudy (anonymous):

(m1*%1+M2%2)=48 m1+m2=100 solve for either m1 or m2 then put it back in eq(1)

OpenStudy (anonymous):

(20m1+55m2)/(m1+m2)=48 (20m1+55m2)/100=48 20m1+55m2=4800 m1+m2=100 m1=100-m2 20(100-m2)+55m2=4800 2000-20m2+55m2=4800 35m2=4800-2000=4600 m2=4600/35=131.428

OpenStudy (perl):

isnt this a system of equations A metalworker has a metal alloy that is 20% copper and another alloy that is 55% copper. How many kilograms of each alloy should the metalworker combine to create 100kg of a 48% copper alloy? x + y = 100 .20x + .55y = .48 * 100

OpenStudy (perl):

where x = # kg of 20% copper alloy y = # kg of 55% copper alloy

OpenStudy (perl):

according to wolfram x = 20, y = 80 http://www.wolframalpha.com/input/?i=solve++x+%2B+y+%3D+100%2C++.20x+%2B+.55y+%3D+.48+*+100+